# MotherNet: A Foundational Hypernetwork for Tabular Classification

### v1

@@ 3.3 MotherNet: Generating Model Weights
Motivated by the success of TabPFN, and inspired by previous work on hypernetworks, we propose MotherNet, a transformer architecture that is trained to produce machine learning models with trained weights in a single forward pass. This methodology combines the benefits of a Foundation Model that does not require dataset specific training or tuning with the high efficiency of a compact model at inference time. The resulting models are small feed-forward neural networks (an MLP with two hidden layers of size 512 in our experiments), that have competitive performance, created without the use of back-propagation or any loss minimization on the training set. The training process of the overall architecture can be described as:
@@
$\begin{split}\min_{\theta}\sum_{i}\mathcal{L}(\texttt{MLP}_{\phi},D^{p}_{i}),\\ \textrm{ where }\phi=\texttt{MotherNet}(\mathbf{D^{t}_{i}},\theta))\end{split}$ (1)
@@
Where $\theta$ are the parameters of the MotherNet transformer, $D^{t}_{i}$ and $D^{p}_{i}$ are training and prediction portion of synthetic dataset $i$, $\textrm{MLP}_{\phi}$ is the feed-forward neural network with parameters given by $\phi$ and $\mathcal{L}(M,D)$ is the cross entropy loss of the model $M$ evaluated on datasets $D$. Training is performed by back-propagation through the whole architecture (from the output of the child model, and through the transformer layers) where each training sample corresponds to one dataset. We will refer to this process as meta-training in the following. During meta-training, the parameters $\theta$ are learned using synthetic datasets, using the prior from Hollmann et al. (2022) and then frozen (as common in Foundation Models). To apply the model to a new (real) dataset $\hat{D}$ consisting of a training portion $\hat{D}^{t}$ and a prediction portion $\hat{D}^{p}$, we evaluate $\texttt{MotherNet}(\hat{D}^{t},\theta)$, which produces a vector of parameters $\hat{\phi}$. This vector is then used as the weight and bias vectors of a feed-forward neural network (properly reshaped), which can be used to make predictions on $\hat{D}^{p}$. We refer to this approach of applying MotherNet to create a child network as in-context learning. This approach is somewhat similar to the approach of Zintgraf et al. (2019), however instead of producing initializations that have to be fine-tuned on a specific dataset, we directly produce the final weights of the model. The absence of gradient descent in our method not only provides an advantage in terms of runtime complexity, it also eliminates the need to apply any regularization, as the model was trained for generalization, not training set performance.
Figure 3: Comparison of TabPFN, MLP-distill and MotherNet with tuned baselines on the test datasets. Left: Mean normalized ROC AUC over time. The two versions of TabPFN overlap too closely to be distinguished. Right: Mean rank over time. Predictions of transformer-based models are constant over the $x$ axis. Compare Figure 9 for validation set results. All transformer based methods use an ensemble of of size 3.
Figure 4: Comparison of normalized mean ROC AUC on test datasets, based on tuning baselines for 1h.
The success of our approach might inspire some skepticism, but we believe our empirical results show not only the feasibility but the competitiveness of the method.
An overview of the model architecture is shown in Figure 1. MotherNet maintains the structure of input encoding and twelve transformer layers of TabPFN on the training set, which produces activations of size $m$ (512 for the experiments) for each pair of training data point and label. On top of this, another attention layer is added, that, based on a learned query vector, reduces all activations to a single dataset embedding $E$ of size $m_{\textrm{all}}$ (2048 in the experiments). The query token is learned during gradient descent and then fixed, similar to prefix tuning (Li & Liang, 2021), but present only in an attention module on top of the transformer layers, not in every layer of the network.
This embedding $E$ is then decoded into the vector $\phi$ using a one-hidden-layer feed-forward neural network. The vector $\phi$ of activations based on the transformer is then reshaped into weights and biases for another feed-forward network.
Initial experiments showed this approach to be very successful, but lead to very large transformer models as a function of the size of the network that was to be produced. Therefore, we employed a standard compression technique and imposed a low rank structure on the weights of the produces networks. As in Ha et al. (2017) we decompose the weights into two components, $W_{i}^{\text{pred}}$ that is the output of the transformer, and $W_{i}^{\text{fixed}}$ that is learned during the meta-training phase and fixed during in-context learning.

### v2

@@ 3.1 MotherNet: Generating Model Weights
Motivated by the success of TabPFN, and inspired by previous work on hypernetworks, we propose MotherNet, a transformer architecture that is trained to produce machine learning models with trained weights in a single forward pass. This methodology combines the benefits of a Foundation Model that does not require dataset specific training or tuning with the high efficiency of a compact model at inference time. The resulting models are small feed-forward neural networks (an MLP with two hidden layers of size 512 in our experiments) that have competitive performance, created without the use of back-propagation or any loss minimization on the training set. The training process of the overall architecture can be described as:
@@
$\begin{split}\min_{\theta}\sum_{i}\mathcal{L}(\texttt{MLP}_{\phi},D^{p}_{i}),\\ \textrm{ where }\phi=\texttt{MotherNet}(D^{t}_{i},\theta))\end{split}$ (1)
@@
Where $\theta$ are the parameters of the MotherNet transformer, $D^{t}_{i}$ and $D^{p}_{i}$ are training and prediction portion of a dataset $i$, $\textrm{MLP}_{\phi}$ is the feed-forward neural network with parameters given by $\phi$ and $\mathcal{L}(M,D)$ is the cross entropy loss of the model $M$ evaluated on datasets $D$. The model architecture is shown in Figure 1. Training is performed by back-propagation through the whole architecture (from the output of the child model, and through the transformer layers) where each training sample corresponds to one dataset. During this meta-training, the parameters $\theta$ are learned using synthetic datasets, using the prior from Hollmann et al. (2022) and then frozen. To apply the model to a new (real) dataset $\hat{D}$ consisting of a training portion $\hat{D}^{t}$ and a prediction portion $\hat{D}^{p}$, we evaluate $\texttt{MotherNet}(\hat{D}^{t},\theta)$, which produces a vector of parameters $\hat{\phi}$. This vector is then used as the weight and bias vectors of a feed-forward neural network (properly reshaped), which can be used to make predictions on $\hat{D}^{p}$. We refer to this approach of applying MotherNet to create a child network as in-context learning. The absence of per-dataset gradient descent in our method not only provides an advantage in terms of runtime complexity, it also eliminates the need to apply and tune regularization, as the model was trained directly for generalization, similar to a Conditional Neural Process (Garnelo et al., 2018).
Figure 2: Comparison of TabPFN, HyperFast, MLP-distill and MotherNet with tuned baselines over the test datasets of Hollmann et al. (2022), listed in Table 8.  † means 1h of HPO on CPU, ‡ means 1h of HPO on GPU. Left: Comparison of normalized mean ROC AUC. Right: Critical Differences Diagram (Demšar, 2006).
MotherNet maintains the structure of input encoding and twelve transformer layers of TabPFN on the training set, which produces embeddings of size $m$ (512 for the experiments) for each pair of training data point and label. We use a one-hot encoding of classes, unlike Hollmann et al. (2022) who use a linear layer. We use the training labels to compute the average embedding per class, reducing all activations to a single dataset embedding $E$ of size $m_{\textrm{all}}$ ($512\cdot 10$ in the experiments). This embedding $E$ is decoded into the vector $\phi$ using a one-hidden-layer feed-forward neural network. The vector $\phi$ of activations based on the transformer is then reshaped into weights and biases for the “child” feed-forward network. We evaluated different variants of the architecture hyper-parameters on the validation set of Hollmann et al. (2022). Initial experiments showed this approach to be very successful, but lead to very large transformer models as a function of the size of the network that was to be produced. To reduce model size, we decomposed the weights into two components, $\mathbf{W}^{\text{p}}_{i}$ that is produced as a prediction by the transformer, and $\mathbf{W}^{\text{f}}_{i}$ that is learned during the meta-training phase and fixed during in-context learning, similar to Ha et al. (2017).